OpenStudy (anonymous):

Determine the limit of (tsint)/(1-cost). I get tripped up as soon as any trig gets involved :)

OpenStudy (gorv):

cos2a=1-2sin^2a take a=t/2 cost=1-2sin^2(t/2) 2sin^2(t/2)=1-cost

OpenStudy (gorv):

tsint/(2sin^2(t/2))

OpenStudy (gorv):

as t tends to zero sint/t=1 use this and solve

OpenStudy (anonymous):

Awesome thank you!

OpenStudy (anonymous):

Ah I forgot to mention, as t approaches 0